coutneythornton:

A dilation with a scale factor of 3 takes triangle PQR to triangle XYZ. What is the ratio of the perimeter of triangle PQR to the perimeter of triangle XYZ?

mxddi3:

do you have a picture of this?

Mercury:

You can solve this without a diagram. notice what happens when you take a generic triangle with side lengths a, b, c and scale them up by some scale factor k. old perimeter = a + b + c new perimeter = ka + kb + kc = k(a+b+c) so the perimeter scales up by k as well apply this logic to your problem